OpenStudy (anonymous):

y^19 * y^0 multiply and simplify. Assume variables represent nonzero real numbers.

OpenStudy (amistre64):

....y^19

OpenStudy (amistre64):

y^0 = 1

OpenStudy (dumbcow):

y^0 = 1 always

OpenStudy (amistre64):

\(y^19 = \sqrt[1/19]{y}\)

OpenStudy (amistre64):

that didnt work like planned lol

myininaya (myininaya):

y^0=1 always?

myininaya (myininaya):

y^0=1 assuming y does not equal 0